OpenStudy (anonymous):

find the points of inflection and discuss the concavity of the graph f(x)=x(x-4)^3

OpenStudy (anonymous):

just having issues with derriving

hartnn (hartnn):

still need help ?

OpenStudy (anonymous):

yes pleaseee :)

hartnn (hartnn):

f(x)=x(x-4)^3 u need to use product rule , tried it ?

OpenStudy (anonymous):

yeah. I just realised that ive been doing the product rule but dividing by g(x)^2 (/.\) EMBERASSINggg x)

hartnn (hartnn):

whats g(x)^2 ??

OpenStudy (anonymous):

I combined it with quotient rule. x)

hartnn (hartnn):

here u only need product rule (fg)' = fg'+f'g f= x, g = (x-4)^3

OpenStudy (anonymous):

im not sure if im deriving right. is the derivative of (x-4)^3,, 3(x-4)^2?

hartnn (hartnn):

yes

hartnn (hartnn):

thats correct, go ahead

OpenStudy (anonymous):

ok. for f''(x)=3x^2-12x+16 (I factored out a 3)

hartnn (hartnn):

lets check, (fg)' = fg'+f'g f= x, g = (x-4)^3 (x (x-4)^3)' = (x-4)^3 + 3x(x-4)^2 = (x-4)^2 (x-4+3x) = 4(x-4)^2 (x-1) did u differentiate this to take 2nd derivative ?

OpenStudy (anonymous):

I derived (x-4)^3+3x(x-4)^2 how did u get to the bottom???

hartnn (hartnn):

i factored out (x-4)^2 , which was there in both terms (x-4)^3 = (x-4)^2 (x-4)

OpenStudy (anonymous):

ohhhhh! Ok

OpenStudy (anonymous):

thanks. I can do the rest from here. thanks you!

hartnn (hartnn):

welcome ^_^